1. 12x-(4x+10)=54
2. you estimate that you spend $115 on groceries each month. how much money do you spend on groceries each week?
3. 8x-10(3-x)=42
4. 2/3(x+4)=8

Answer by rfer(16322)

1)
12x-(4x+10)=54
12x-4x-10=54
8x=64
x=8
2)
115/4=$28.75 per wk
3)
8x-10(3-x)=42
8x-30+10x=42
18x=72
x=4
4)
you try this one